A random number generator picks a number from 1 to 21 in a uniform manner. Round all answers to two decimal places.

A. The mean of this distribution is 11

B. The standard deviation is 5.77

C. The probability that the number will be exactly 15 is P(x = 15) = 0

D. The probability that the number will be between 4 and 8 is P(4 < x < 8) = 0.20

E. The probability that the number will be larger than 15 is P(x > 15) = 0.30

F. P(x > 5 | x < 11) =

G. Find the 60th percentile.

H. Find the minimum for the upper quartile.

I don't have the answers to F.G.H thanks a lot for your help by the way
Answer #1

F

P(X>5 |X<11) =P(5<X<11)/P(X<11) =(11-5)/(11-1)=0.6

G)

pth percentile =a+p*(b-a)/100   
60th percentile =1+60*(21-1)/100=13

H) minimum for the upper quartile(75th percentile) =1+75*(21-1) =16